OpenStudy (anonymous):

How do you solve for 'h' in the equation V=πr^2h

OpenStudy (anonymous):

You need to get h all by itself on one side of the equation. You could start by dividing both sides by pi, which will remove the pi from the right side

OpenStudy (jdoe0001):

\(\bf V = \pi r^2 h \implies \cfrac{V}{ \pi r^2} = h\)

OpenStudy (jdoe0001):

as suggested, divide both sides by pi, then divide again both sides by r^2